# Aggregated Attributions for Explanatory Analysis of 3D Segmentation Models

### v2

@@ Abstract
Analysis of 3D segmentation models, especially in the context of medical imaging, is often limited to segmentation performance metrics that overlook the crucial aspect of explainability and bias. Currently, effectively explaining these models with saliency maps is challenging due to the high dimensions of input images multiplied by the ever-growing number of segmented class labels. To this end, we introduce Agg2Exp, a methodology for aggregating fine-grained voxel attributions of the segmentation model’s predictions. Unlike classical explanation methods that primarily focus on the local feature attribution, Agg2Exp enables a more comprehensive global view on the importance of predicted segments in 3D images. Our benchmarking experiments show that gradient-based voxel attributions are more faithful to the model’s predictions than perturbation-based explanations. As a concrete use-case, we apply Agg2Exp to discover knowledge acquired by the Swin UNEt TRansformer model trained on the TotalSegmentator v2 dataset for segmenting anatomical structures in computed tomography medical images. Agg2Exp facilitates the explanatory analysis of large segmentation models beyond their predictive performance.
@@ 1 Introduction
Semantic segmentation of 3D images is a crucial task in computer vision, with useful applications in domains such as robotics [11], autonomous vehicles [57], augmented reality [21], and healthcare [41, 27, 24, 17]. Notably, an automated and accurate segmentation of 3D medical images like computed tomography (CT) scans has the potential to significantly improve the experiences of both clinicians and patients [51, 20, 48, 17]. The motivation to aid humans in overcoming diseases [40, 18] has lead to continuous improvements of 3D segmentation models [32, 41, 27, 33, 24, 23, 64, 49, 36, 70, 10], which are trained on an ever-increasing amount of medical data [47, 46, 34, 3, 64, 2].
Large and complex 3D segmentation models become black-box predictive systems, but they were first intended to collaborate with humans. Unavoidably, the ability to provide understandable, or explainable, model outputs is limited and often leads to trust issues [14, 37, 61]. In extreme cases, the lack of understandability conceals biases in data, which negatively influence the model’s generalization performance [30, 38], or even make it discriminate certain groups of patients [19].
To this end, various post-hoc explanation methods have been proposed that aim to approximate the machine learning model’s reasoning [61, 69, 5, 38]; the most popular class of methods being feature attributions (aka saliency maps[58, 52, 45, 59, 56, 63, 67, 8, 35, 1]. The quality of feature attribution methods can be benchmarked with evaluation metrics [4, 28, 39], including faithfulness to the model [66, 65] and sensitivity to data perturbations [68, 6], which facilitates the work on their improvement. Recently, researchers proposed pixel attribution methods specific to 2D segmentation models including gradient-based attributions [29, 12], model-agnostic perturbation-based attributions [29, 15], as well as class activation mapping [62, 22] and layer-wise relevance propagation [16] that are specific to convolutional neural networks.
@@
    We introduce the Aggregate2Explain (Agg2Exp) methodology for aggregating attributions into segment importance explanations. Agg2Exp overcomes the inherit limitations of local feature attributions, like high data dimensionality and model complexity, which are specific to the task of 3D image segmentation.
@@
    As a concrete use-case, we apply Agg2Exp to analyze the state-of-the-art model for segmenting anatomical structures in 3D CT scans. Agg2Exp allows discovering model biases and explaining its abnormal behavior.
@@ 2.2 Explanation of (2D) segmentation models
Unlike for image classification [14, 61, 37], only a few works consider explaining 2D segmentation models [61, 69, 5, 38]. The most popular approach we focus on is to measure the feature (pixel) importance and visualize it as an attribution (saliency) map [52, 59, 56, 63]. In [29], the first pixel attribution methods for semantic segmentation are introduced in the form of a saliency grid approximated with gradient-based or perturbation-based algorithms. Follow-up work considers extending gradient-weighted class activation mapping to the segmentation task [62, 22]. In [15], the authors consider a Shapley-based method that relies on developing an interpretable image representation, e.g. with super-pixels (similar to LIME [52]). It is noteworthy that pixel-wise attributions can be used to improve segmentation models [12, 16]. More broadly related works consider counterfactual explanations to analyze prediction errors [69], concept-based explanations to discover bias [16], and inherently interpretable deep neural networks to overcome the issue of explaining models post-hoc [55, 25].
None of the above mentioned works has considered explaining 3D segmentation models, which poses significant algorithmic, computational and perceptive challenges as the data dimension increases exponentially both in the input and output space. We propose an aggregation framework to fill this research gap where explanations of 3D segmentation are visually more comprehensible and thus provide added value in the analysis of medical imaging models specific to 3D segmentation.
@@ 3.1 Voxel attribution methods for 3D segmentation
Integrated gradients (IG) aggregates VG on a path from $\mathbf{x}$ to the selected baseline input $\mathbf{x}^{\prime}$ (usually $\mathbf{x}^{\prime}=\mathbf{0}$):
@@
Perturbation-based attribution methods are model-agnostic [52], i.e. they treat a model as a black-box without assuming any knowledge about its structure unlike gradient-based methods. We adapt KernelSHAP [45, 15] to explain $\hat{f}_{c}$, which requires defining an interpretable feature representation. In case of 3D images, we propose the following two supervoxel representations: (1) Semantic: Given a prediction $\hat{\mathbf{y}}$ for an input $\mathbf{x}$, its semantic supervoxel partitioning is defined as $\mathbf{v}^{(i)}\coloneq\operatorname*{argmax}_{j}\hat{\mathbf{y}}^{(i,j)}$, where $\mathbf{v}\in[1,\ldots,l]^{p}$. (2) Cubes: Given a predefined number of regions $r$, an input $\mathbf{x}$ is partitioned into supervoxels $\mathbf{v}\in[1,\ldots,r]^{p}$ defined by even cubes, e.g. an image of size $512^{3}$ can be partitioned into $64$ regions, each of size $128^{3}$.
@@ 3.2 Aggregating attributions for 3D segmentation
Surprisingly, our paper is the first to point out that voxel attribution explanations of 3D segmentation become increasingly complex and impractical (cf. Fig. 2). For example, given $512^{3}$ voxels and 17 class labels, one obtains about $10^{9}$ voxel attributions explaining a single model prediction. To overcome this challenge of high complexity stemming from the multiplication of input and output dimensions, we propose to aggregate voxel attributions into simpler explanations that are easier to visualize and analyze.
@@
    Local RoI importance. Aggregate voxel attributions based on a predefined set of RoIs, which include the segmented class labels. This step lowers the complexity of explanations. Note that the set can be extended by other RoIs independent of the model’s predictions, e.g. an output from another segmentation model or annotated by a human.
@@
    Global RoI importance. Aggregate local RoI importances with an average over a set of inputs to obtain global importance. This step enables a more comprehensive global view on the 3D segmentation model.
@@
where the correlation is estimated with bootstrapping over $n$ rounds of sampling a subset $S_{i}$ of size $m$. We desire an attribution method $g$ to have high faithfulness.
@@ 4.1 Setup
To evaluate Agg2Exp, we use two datasets: the publicly available TotalSegmentator v2 [TSV2, 64] test set, and our private dataset (denoted as B50), [^1] An ethical approval was granted by the Ethics Committee of the Medical University of Warsaw. comprising 32 and 127 CT scans, respectively. The inclusion of the latter is motivated by our real-world application of Agg2Exp in practice to analyze the 3D segmentation model that will potentially be deployed in a hospital software. Both datasets consist of 3D CT scans of humans’ chest, which include healthy (aka normal) cases, as well as cases with various thoracic pathologies indicating illness. Following [64], we select 16 of the most important anatomical structures of the thorax as the class labels for segmentation, which, including background, results in a set of 17 labels in total. Note that the above mentioned CT scans have all those classes present. As a preprocessing step, we standardize CT scans to an isotropic resolution of $1.5^{3}\;\mathrm{mm}^{3}$, clip Hounsfield Units into $[-1024,1024]$ and then scale them linearly to the range $[0,1]$.
@@
We use the well-established Swin UNETR [23, 60] architecture as a baseline model for our experiments. It is a faster [^2] Model inference performs about 10×10$\times$10 × faster due to changes in model size and architecture. and more accessible [^3] TotalSegmentator is closed behind a complex Python API, which, for example, prevents its accessible differentiation to approximate gradients. version of the original TotalSegmentator model [64] that is based on the older nnUNet [32] architecture. We train the model on the TSV2 train set. We observe performance improvements specifically due to merging multiple ribs and vertebraes into a single class, as well as carefully selecting the 16 anatomical structures. Overall, it lowers the computational complexity while allowing us to segment anatomical structures vital for our collaborating radiologists.
@@
Both the private dataset B50 and the network weights of our trained model are available upon a reasonable request. We provide further details on the hyperparameters of the voxel attribution methods and their evaluation metrics in Appendix B. Code to reproduce the experimental results will be available on GitHub at https://github.com/MI2DataLab/agg2exp.
@@ 4.2 Benchmarking voxel attribution methods
Table 1: Benchmark of 4 voxel attribution methods (including KernelSHAP in the two variants) with 4 evaluation metrics on 2 validation datasets. We report averages and standard deviations over a set of inputs and class labels, and underline the best scores in each metric.
@@
We benchmark voxel attribution methods introduced in Sec. 3.1 with evaluation metrics introduced in Sec. 3.3 in our setup. Table 1 shows results where we report an average of the metrics’ values over all patient cases and seven class labels, [^4] Class labels: background, aorta, lung lower lobe left, lung lower lobe right, trachea, heart, ribs. which is due to high computational complexity of calculating those metrics, i.e. this benchmark took around two weeks to compute on an 8$\times$×A100 GPU node, highlighting the particular complexity of this task. KernelSHAP is estimated based on the two introduced supervoxel representations, where semantic considers all 17 class labels, and cubes is based on about $64$ regions, each of size $128^{3}$, depending on the size of a scan. In general, the two KernelSHAP variants are significantly less efficient than gradient-based methods because they need to perform a large number of model inference steps. Gradient-based attribution methods have lower complexity than the perturbation-based KernelSHAP because the latter attributes larger parts of the input as important.
@@ 4.3 Use-case: Explanatory analysis of a model for segmenting anatomical structures in CT scans
Fig. 3 shows the distribution of local RoI importances for two exemplary class labels: ribs and pulmonary vein. Following [64], we color objects by their semantic meaning, i.e., cardiovascular system in red, muscles in yellow, bones in blue, other organs in grey, and lung pathologies in black. Masks for the latter are available in our B50 dataset, which allows us to extend the analysis to regions that are of interest for the segmentation of chest CT scans. The model segments ribs relying mostly on the part of input classified as ribs itself. Contrary, segmenting a pulmonary vein requires the model to use a broader context, i.e. the regions of heart and lungs highly contribute to pulmonary vein segmentation. Crucially, for many input cases (outliers denoted with black dots), regions of the consolidation and pleural effusion pathologies influence the model’s predictions.
@@
We use Agg2Exp to perform local outlier analysis, i.e. find patient cases for which the model behaves abnormally. We apply the rather straightforward Isolation Forest [IF, 42] algorithm to the matrix consisting of local RoI importances. We do that for each of the 17 class labels separately, resulting in an IF trained on 17 features from Agg2Exp per each class label. We first train IF on the TSV2 training set in order to model the normal behavior and then apply it on validation data (in this case our B50 dataset).
@@
Additionally, we analyze whether the abnormal model’s behavior entails worse predictive performance (cf.Sec. C.1). We conduct the Spearman’s rank test between the anomaly score returned by IFs and the Dice score from TSV2 test data for each class label, which shows a statistically significant correlation for some labels (including lungs). It further suggests that the aggregated attributions can highlight weak model performance for specific labels in inference time, i.e. without the ground truth segmentation masks.
We also conduct a Spearman’s rank correlation test to analyze the relationship between anomaly scores averaged over labels and the average Dice scores. The test returns a p-value of $\approx 0.0024$, which shows a statistically significant relation between the average anomaly score and the average Dice score.
@@ 5 Limitations and future work
We identify three main limitations of Agg2Exp that provide natural opportunities for future work. By design, Agg2Exp inherits the general limitations of feature attribution methods [54, 37], which at best approximate an explanation of the model’s predictions. As a countermeasure, we proposed to perform a benchmark considering several desiderata to guide a choice of the most suitable voxel attribution method for a given applied use-case. Certainly, improvements in the faithfulness and efficiency of voxel attribution approximation will lead to improvements of Agg2Exp [16, 66]. Second, our proposition of using an aggregated model function (Eq. 2) as a proxy for an explanation could be suboptimal. Although it was proven sufficient for our use-case, future work can consider alternative aggregation measures to achieve different goals. Finally, we acknowledge that Agg2Exp relates to human-computer interaction in that explanations can be misunderstood by users or even manipulate their perception [50]. Following the arguments posited in [7], we believe Agg2Exp to answer the needs of machine learning developers and radiologists debugging our segmentation model, leaving an evaluation via user studies as a future research direction [53].

### v3

@@ Abstract
Analysis of 3D segmentation models, especially in the context of medical imaging, is often limited to segmentation performance metrics that overlook the crucial aspect of explainability and bias. Currently, effectively explaining these models with saliency maps is challenging due to the high dimensions of input images multiplied by the ever-growing number of segmented class labels. To this end, we introduce Agg2Exp, a methodology for aggregating fine-grained voxel attributions of the segmentation model’s predictions. Unlike classical explanation methods that primarily focus on the local feature attribution, Agg2Exp enables a more comprehensive global view on the importance of predicted segments in 3D images. Our benchmarking experiments show that gradient-based voxel attributions are more faithful to the model’s predictions than perturbation-based explanations. As a concrete use-case, we apply Agg2Exp to discover knowledge acquired by the Swin UNEt TRansformer model trained on the TotalSegmentator v2 dataset for segmenting anatomical structures in computed tomography medical images. Agg2Exp facilitates the explanatory analysis of large segmentation models beyond their predictive performance. The source code is publicly available at https://github.com/mi2datalab/agg2exp.
@@ 1 Introduction
Semantic segmentation of 3D images is a crucial task in computer vision, with useful applications in domains such as robotics [12], autonomous vehicles [59], augmented reality [21], and healthcare [42, 27, 24, 50, 18, 63]. Notably, automated and accurate segmentation of 3D medical images like computed tomography (CT) scans has the potential to significantly improve the experiences of both clinicians and patients [52, 66, 48, 18]. The motivation to aid humans in overcoming diseases [41, 19] has lead to continuous improvements of 3D segmentation models [32, 42, 27, 33, 24, 23, 69, 49, 37, 75, 11], which are trained on an ever-increasing amount of medical data [47, 46, 35, 3, 69, 2].
Large and complex 3D segmentation models become black-box predictive systems, which were initially developed to aid human decision-making. Unavoidably, the ability to provide understandable, or explainable, model outputs is limited and often leads to trust issues [15, 38, 65]. In extreme cases, the lack of understandability conceals biases in data, which negatively influence the model’s generalization performance [30, 39], or even make it discriminate certain groups of patients [20].
To this end, various post-hoc explanation methods have been proposed that aim to approximate the machine learning model’s reasoning [65, 74, 5, 39, 61]; the most popular class of methods being feature attributions [60, 53, 57, 62, 58, 68, 72, 9, 36, 1]. The quality of feature attribution methods can be benchmarked with evaluation metrics [4, 28, 40], including faithfulness to the model [71, 70] and sensitivity to data perturbations [73, 7], which facilitates the work on their improvement. Recently, researchers proposed pixel attribution methods specific to 2D segmentation models including gradient-based attributions [29, 13], model-agnostic perturbation-based attributions [29, 16], as well as class activation mapping [67, 22] and layer-wise relevance propagation [17] that are specific to convolutional neural networks.
@@
    We introduce the Aggregate2Explain (Agg2Exp) methodology for aggregating attributions into segment importance explanations. Agg2Exp overcomes the inherited limitations of local feature attributions, like high data dimensionality and model complexity, which are specific to the task of 3D image segmentation.
@@
    As a concrete use-case, we apply Agg2Exp to analyze the state-of-the-art model for segmenting anatomical structures in 3D CT scans. Agg2Exp allows discovering model biases and explaining their abnormal behavior.
@@ 2.2 Explanation of (2D) segmentation models
Unlike for image classification [15, 65, 38], only a few works consider explaining 2D segmentation models [65, 74, 5, 39]. The most popular approach we focus on is to measure the feature (pixel) importance and visualize it as an attribution (saliency) map [53, 62, 58, 68]. In [29], the first-pixel attribution methods for semantic segmentation are introduced in the form of a saliency grid approximated with gradient-based or perturbation-based algorithms. Follow-up work considers extending gradient-weighted class activation mapping to the segmentation task [67, 22]. In [16], the authors consider a Shapley-based method that relies on developing an interpretable image representation, e.g. with super-pixels (similar to LIME [53]). It is noteworthy that pixel-wise attributions can be used to improve segmentation models [13, 17]. More broadly related works consider counterfactual explanations to analyze prediction errors [74], concept-based explanations to discover bias [17], and red-teaming analysis [34]. Another approach is training inherently interpretable neural networks to overcome the issue of explaining models post-hoc [56, 25].
None of the above-mentioned works has considered explaining 3D segmentation models, which pose significant algorithmic, computational, and perceptive challenges as the data dimension increases exponentially both in the input and output space. We propose an aggregation framework to fill this research gap where explanations of 3D segmentation are visually more comprehensible and thus provide added value in the analysis of medical imaging models specific to 3D segmentation.
@@ 3.1 Voxel attribution methods for 3D segmentation
Integrated gradients (IG) aggregate VG on a path from $\mathbf{x}$ to the selected baseline input $\mathbf{x}^{\prime}$ (usually $\mathbf{x}^{\prime}=\mathbf{0}$):
@@
Perturbation-based attribution methods are model-agnostic [53], i.e. they treat a model as a black-box without assuming any knowledge about its structure, unlike gradient-based methods. We adapt KernelSHAP [57, 16] to explain $\hat{f}_{c}$, which requires defining an interpretable feature representation. In case of 3D images, we propose the following two supervoxel representations: (1) Semantic: Given a prediction $\hat{\mathbf{y}}$ for an input $\mathbf{x}$, its semantic supervoxel partitioning is defined as $\mathbf{v}^{(i)}\coloneq\operatorname*{argmax}_{j}\hat{\mathbf{y}}^{(i,j)}$, where $\mathbf{v}\in[1,\ldots,l]^{p}$. (2) Cubes: Given a predefined number of regions $r$, an input $\mathbf{x}$ is partitioned into supervoxels $\mathbf{v}\in[1,\ldots,r]^{p}$ defined by even cubes, e.g. an image of size $512^{3}$ can be partitioned into $64$ regions, each of size $128^{3}$.
@@ 3.2 Aggregating attributions for 3D segmentation
Surprisingly, our paper is the first to point out that voxel attribution explanations of 3D segmentation have become increasingly complex and impractical (cf. Fig. 2). For example, given $512^{3}$ voxels and 17 class labels, one obtains about $10^{9}$ voxel attributions explaining a single model prediction. To overcome this challenge of high complexity stemming from the multiplication of input and output dimensions, we propose to aggregate voxel attributions into simpler explanations that are easier to visualize and analyze.
@@
    Local RoI importance. Aggregate voxel attributions based on a predefined set of RoIs, which include the segmented class labels. This step lowers the complexity of explanations. Note that the set can be extended by other RoIs independent of the model’s predictions, e.g. output from another segmentation model or annotated by a human.
@@
    Global RoI importance. Aggregate local RoI importance with an average over a set of inputs to obtain global importance. This step enables a more comprehensive global view of the 3D segmentation model.
@@
where the correlation is estimated with bootstrapping over $n$ rounds of sampling a subset $S_{i}$ of size $m$. We desire an attribution method $g$ to be highly faithful.
@@ 4.1 Setup
To evaluate Agg2Exp, we use two datasets: the publicly available TotalSegmentator v2 [TSV2, 69] test set, and our private dataset (denoted as B50), [^1] An ethical approval was granted by the Ethics Committee of the Medical University of Warsaw. comprising 32 and 127 CT scans, respectively. The inclusion of the latter is motivated by our real-world application of Agg2Exp in practice to analyze the 3D segmentation model that will potentially be deployed in hospital software. Both datasets consist of 3D CT scans of humans’ chests, which include healthy (normal) cases, as well as cases with various thoracic pathologies indicating illness. Following [69], we select 16 of the most important anatomical structures of the thorax as the class labels for segmentation, which, including background, results in a set of 17 labels in total. Note that the above-mentioned CT scans have all those classes present. As a preprocessing step, we standardize CT scans to an isotropic resolution of 1.5 $\times$ 1.5 $\times$ 1.5 $\mathrm{mm}^{3}$, clip Hounsfield units to the range $[-1024,1024]$, and then scale them linearly to $[0,1]$.
@@
We use the well-established Swin UNETR [23, 64] architecture as a baseline model for our experiments. It is a faster [^2] Model inference performs about 10×10$\times$10 × faster due to changes in model size and architecture. and more accessible [^3] TotalSegmentator is closed behind a complex Python API, which, for example, prevents its accessible differentiation to approximate gradients. version of the original TotalSegmentator model [69] that is based on the older nnUNet [32] architecture. We train the model on the TSV2 train set. We observe performance improvements specifically due to merging multiple ribs and vertebrae into a single class, as well as carefully selecting the 16 anatomical structures. Overall, it lowers the computational complexity while allowing us to segment anatomical structures vital for our collaborating radiologists.
@@
Both the private dataset B50 and the network weights of our trained model are available upon a reasonable request. We provide further details on the hyperparameters of the voxel attribution methods and their evaluation metrics in Appendix B.
@@ 4.2 Benchmarking voxel attribution methods
Table 1: Benchmark of 4 voxel attribution methods (including KernelSHAP in the two variants) with 4 evaluation metrics on 2 validation datasets. We report means and standard deviations over a set of inputs and class labels, and underline the best scores in each metric.
@@
We benchmark voxel attribution methods introduced in Sec. 3.1 with evaluation metrics introduced in Sec. 3.3 in our setup. Table 1 shows results where we report a mean of the metrics’ values over all patient cases and seven class labels, [^4] Class labels: background, aorta, lung lower lobe left, lung lower lobe right, trachea, heart, ribs. which is due to high computational complexity of calculating those metrics, i.e. this benchmark took around two weeks to compute on an 8$\times$×A100 GPU node, highlighting the particular complexity of this task. KernelSHAP is estimated based on the two introduced supervoxel representations, where semantic considers all 17 class labels, and cubes is based on about $64$ regions, each of size $128^{3}$, depending on the size of a scan. In general, the two KernelSHAP variants are significantly less efficient than gradient-based methods because they need to perform a large number of model inference steps. Gradient-based attribution methods have lower complexity than the perturbation-based KernelSHAP because the latter attributes larger parts of the input as important.
@@ 4.3 Use-case: Explanatory analysis of a model for segmenting anatomical structures in CT scans
Fig. 3 shows the distribution of local RoI importances for two exemplary class labels: ribs and pulmonary vein. Following [69], we color objects by their semantic meaning, i.e., the cardiovascular system in red, muscles in yellow, bones in blue, other organs in grey, and lung pathologies in black. Masks for the latter are available in our B50 dataset, which allows us to extend the analysis to regions that are of interest for the segmentation of chest CT scans. The model segments ribs by relying mostly on the part of input classified as ribs itself. Contrary, segmenting a pulmonary vein requires the model to use a broader context, i.e. the regions of heart and lungs highly contribute to pulmonary vein segmentation. Crucially, for many input cases (outliers denoted with black dots), regions of the consolidation and pleural effusion pathologies influence the model’s predictions.
@@
We use Agg2Exp to perform local outlier analysis, i.e. find patient cases for which the model behaves abnormally. We apply the rather straightforward Isolation Forest [IF, 43] algorithm to the matrix consisting of local RoI importances. We do that for each of the 17 class labels separately, resulting in an IF trained on 17 features from Agg2Exp per each class label. We first train IF on the TSV2 training set in order to model the normal behavior and then apply it to validation data (in this case our B50 dataset).
@@
Additionally, we analyze whether the abnormal model’s behavior entails worse predictive performance (cf.Sec. C.1). We conduct Spearman’s rank test between the anomaly score returned by IFs and the DSC from TSV2 test data for each class label, which shows a statistically significant correlation for some labels (including lungs). It further suggests that the aggregated attributions can highlight weak model performance for specific labels in inference time, i.e. without the ground truth segmentation masks.
We also conduct a Spearman’s rank correlation test to analyze the relationship between anomaly scores averaged over labels and the mean DSC. The test returns a p-value of $\approx 0.0024$, which shows a statistically significant relation between the mean anomaly score and the mean DSC.
@@ 5 Limitations and future work
We identify three main limitations of Agg2Exp that provide natural opportunities for future work. By design, Agg2Exp inherits the general limitations of feature attribution methods [55, 38], which at best approximate an explanation of the model’s predictions. As a countermeasure, we proposed to perform a benchmark considering several desiderata to guide a choice of the most suitable voxel attribution method for a given applied use-case. Certainly, improvements in the faithfulness and efficiency of voxel attribution approximation will lead to improvements of Agg2Exp [17, 71]. Second, our proposition of using an aggregated model function (Eq. 2) as a proxy for an explanation could be suboptimal. Although it was proven sufficient for our use-case, future work can consider alternative aggregation measures to achieve different goals. Finally, we acknowledge that Agg2Exp relates to human-computer interaction in that explanations can be misunderstood by users [6] or even manipulate their perception [51]. Following the arguments posited in [8], we believe Agg2Exp to answer the needs of machine learning developers and radiologists debugging our segmentation model, leaving an evaluation via user studies as a future research direction [54].